OpenStudy (anonymous):

Please Help!! Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (phi):

find the angle and distance (i.e. r) for the point (2,-2) another polar coordinate will be 180º + that angle ( or -180º + that angle) and make r *negative*

OpenStudy (phi):

You can think of a polar coordinate as first turn in the specified direction and then walk "r" steps in that direction. If r is *negative*, it means walk backwards!

OpenStudy (anonymous):

so (2,-2+180) and (2,-2-180)??

OpenStudy (phi):

except (2,-2) are (x,y) coordinates you need to find the "length" (distance from 0,0 to (2,-2) ) that will be r use the "distance formula" and find the angle. arc tangent of -2/2

OpenStudy (anonymous):

Find the angle using tan = y/x......tan angle = 2/2 or a 45° angle in quad 4......(x,y) = (rcos , r sin)

OpenStudy (phi):